OpenStudy (anonymous):

What is the simplified form of (9x^2-25)/(3x+5)? choices: 3x + 5, with the restriction x ≠ 5/3 3x – 5, with the restriction x ≠ 5/3 3x – 5, with the restriction x ≠ – 5/3 3x + 5, with the restriction x ≠ –5/3

OpenStudy (anonymous):

Here 9x^2 goes into 3x 3x times. (3x * 3x) = 9x^2. 5 goes into 25, 5 times. Take the negative with you. What times 3 = -5??

OpenStudy (anonymous):

the third one coz the only cancell with (3x+5)>>3*(-5/3)+5=0

OpenStudy (anonymous):

@qpHalcy0n, nothing equals -5 multiplied by 3 .

OpenStudy (anonymous):

@moha_10 i didnt think of doing the problem this way , thanks (:

OpenStudy (anonymous):

Yes, it does. 3 * (-5/3) = -5.

OpenStudy (anonymous):

ohhh aha i was using a calc and i got a decimal . wowowwww ! im such a ditz lmao thanks (: